# Natural Language Actor-Critic: Scalable Off-Policy Learning in Language Space

### v1

@@ 4.2 Policy Improvement
Thus far, we showed how to train the natural language critic to evaluate a fixed policy $\pi$. We now show how an improved policy can be learned using textual critiques $Q_{L}^{\pi}(s,a)$ obtained by a critic using Equation 3. Naturally, such policy is a greedy policy where $a\sim\pi(\cdot|s)$ satisfies $a=\arg\max_{a^{\prime}}Q^{\pi}_{L}(s,a^{\prime})$. Note that we assume the following:
@@
For any policy $\pi$, the set $\{Q_{L}^{\pi}(s,a^{\prime})\}_{a^{\prime}\in\mathcal{A}}$ computed using Equation 3 for any state $s$ forms a totally-ordered set with binary relation $\geq$.
@@
However, computing the greedy policy is intractable for LLM agent tasks, where the action spaces $\mathcal{A}\subseteq\mathcal{V}^{*}$ are combinatorial in the token vocabulary, making it impossible to enumerate all possible actions to find the optimal one. While prior works have proposed sampling a subset of actions and reweighting (li2024QProbe), we find empirically that for tractable sample sizes, this approach does not sufficiently explore the space of possible actions.
@@
Finally, we propose a policy improvement objective from $\pi$ to $\pi^{\prime}$ that projects the refinement policy back to the base policy, similar to the policy updates in SAC (haarnoja2018soft). However, rather than parameterizing a target policy using the learned values, which requires enumeration over actions and is intractable in our setting, we let the target policy be the refinement policy:
$\pi^{\prime}=\arg\max_{\pi}\mathbb{E}_{s_{t}\sim\mathcal{D}}\left[D_{f}\left(\pi(\cdot\mid s_{t})\ ||\ \pi^{r}(\cdot\mid s_{t},a_{t}^{1},\ldots,Q^{\pi}_{L}(s,a_{t}^{m}))\right)\right]\,.$ (4)
@@
5 Theoretical Analysis
The goal of this section is to show that policy iteration using our proposed NLAC method converges to the optimal policy. Due to space, we only state our main results and defer full proofs to Appendix A. For traditional actor-critic algorithms, this involves showing (1) convergence of learned Q-values via the Bellman backup, and (2) monotonic improvement of the trained policy. However, such analysis does not apply because the Q-values we learn are textual rather than scalar.
Assumption 5.1.
Assumption 5.2.
For any state $s$ and action $a$ , the language Bellman backup satisfies
@@ 5 Theoretical Analysis
Consider policy evaluation via Equation 2 and let $Q_{L}^{\pi}$ be the natural language critic at convergence. For any state $s$ and action $a$, there exists monotonic mapping $g$ such that $Q^{\pi}(s,a)=g(Q_{L}^{\pi}(s,a))$, where $Q^{\pi}$ denotes the true scalar Q-function.
Our main result makes a precise connection between our learned critic and the true Q-function. The crux of our proof involves showing that our training objective results in a fixed point where $\Phi^{\pi}(s,a)$ are successor features of the underlying MDP (barreto2018successorfeaturestransferreinforcement).
During policy improvement, we update the policy towards a refinement policy $\pi^{r}$. By definition, the refinement policy is an oracle that for any state $s$ and action $a$, generates $a^{r}$ such that $Q_{L}(s,a^{r})\geq Q_{L}(s,a)$. Combining this with the result of Theorem 5.1, we arrive at our second main result:
Theorem 5.2.
Repeated application of policy evaluation via Equation 2 and policy improvement via Equation 4 from a policy $\pi_{0}$ converges to policy $\pi^{*}$ such that $Q^{\pi^{*}}(s,a)\geq Q^{\pi}(s,a)$ for any state $s$ and action $a$, and other policy $\pi$.
Hence, we are able to show that under the aforementioned assumptions, our approach NLAC can provably find the optimal policy for an underlying MDP. Next, we show how to approximately implement NLAC in a scalable and practical algorithm.
@@ 6.1 Training Components
Policy. Many prior works have parameterized policies as LLMs. One of the greatest advantages of doing so is the ability to leverage the strong reasoning capabilities of LLMs from chain-of-thought prompting wei2023chainofthought; yao2022react. By choosing a proper prompt $p_{\text{react}}$, an LLM policy can be instructed to describe their underlying thoughts for choosing a particular action in addition to generating the action itself $a_{t}\sim\mathcal{M}_{\theta}(p_{\text{react}}(s_{t}))$.
Language successor model. LLMs have demonstrated efficacy at predicting realistic future rollouts in a variety of environments (lin2024learningmodelworldlanguage). These futures are generated by simply processing the state-action in a prediction prompt $p_{\text{pred}}$ that also instructs the LLM to summarize rollouts into concise textual descriptions, then sampling from the LLM output $M_{\theta}(\cdot\mid s_{t},a_{t})=\mathcal{M}_{\theta}(p_{\text{pred}}(s_{t},a_{t}))$.
@@
$\mathcal{B}_{L}M_{\theta}(\cdot\mid s_{t},a_{t})=\mathcal{M}_{\theta}(p_{\text{tpred}}(r_{t},s_{t+1},d_{t+1}))\,,\quad d_{t+1}\sim\mathcal{M}_{\theta}(p_{\text{pred}}(s_{t+1}))\,.$
@@
$Q_{L}^{\theta}(s_{t},a_{t})=E_{\theta}(s_{t},a_{t},d_{t}^{(1)},\ldots,d_{t}^{(k)})\sim\mathcal{M}_{\theta}(p_{\text{eval}}(d_{t}^{(1)},\ldots,d_{t}^{(k)}))\,.$
Refinement policy. Finally, the refinement policy $\pi^{r}$ can also be obtained by an LLM instructed to refine its latest action given an evaluation similar to prior self-refinement approaches (madaan2023selfrefineiterativerefinementselffeedback). The refined action is obtained via prompt $p_{\text{refine}}$ as $a_{t}^{r}\sim\mathcal{M}_{\theta}(p_{\text{refine}}(s_{t},a_{t}^{1},\ldots,Q^{\theta}_{L}(s_{t},a_{t}^{m})))$.
@@ 6.2 Training Algorithm
$\mathcal{L}_{1}(s_{t},a_{t},r_{t},s_{t+1})=D_{\mathsf{KL}}\left(\mathcal{B}_{L}M_{\bar{\theta}}(\cdot\mid s_{t},a_{t})\ ||\ M_{\theta}(\cdot\mid s_{t},a_{t})\right)\,$, (5)
where $\bar{\theta}$ are reference parameters that are an exponentially moving average of the trained parameters, in order to prevent generative collapse  (Shumailov2024AIMC). We choose the reverse direction of KL-divergence to capture the full diversity over possible futures. Then, for policy improvement, we train the policy on the log-likelihood loss:
$\qquad\qquad\quad\ \mathcal{L}_{2}(s_{t},k)$ $=-\log\pi_{\theta}(a_{t}^{r}\mid s_{t})\,,\quad a_{t}\!\sim\!\pi_{\theta}(\cdot\mid s_{t})\,,\ a_{t}^{r}\!\sim\!\pi^{r}_{\theta}(\cdot\mid s_{t},a_{t},Q_{L}^{\theta}(s_{t},a_{t}))\,.$ (6)
This objective can be interpreted as distillation, but using generations by the refinement policy as the teacher policy. Note that the loss depends on $k$ via $Q_{L}^{\theta}(s_{t},a_{t})$ given by Equation 3. Note that by default, our refinement policy relies on the base reasoning capabilities of the pretrained LLM. In Appendix B.5, we show results when the refinement policy is explicitly trained.
@@
We show pseudocode for NLAC in Algorithm 1. In practice, we found it helpful to implement $\mathcal{D}$ as a prioritized replay buffer weighted by $\mathcal{L}_{1}(s_{t},a_{t},s_{t+1})$ with sampling parameter $\alpha$ (schaul2016prioritizedexperiencereplay). This is because in many tasks, though a base LLM policy may achieve low reward in a large proportion of rollouts, many actions in these unsuccessful rollouts are still optimal. Therefore, to improve learning efficiency, we prioritize training on samples where the agent is likely to take a suboptimal action, using critic loss as a proxy for the likelihood.
Like other methods that utilize pretrained LLMs, our method is susceptible to catastrophic forgetting. We were able to avoid this in our experiments by training in low-data regimes. However, we discuss effective methods for mitigating catastrophic forgetting in Appendix B.6.
@@ 7 Experiments
To demonstrate the effectiveness of NLAC, we evaluate our method on a variety of LLM agent tasks: mathematical reasoning (hendrycksmath2021), strategic dialogue  (20Q), and customer service using mixed dialogue and tool-use (yao2024taubenchbenchmarktoolagentuserinteraction). Though mathematical reasoning does not involve interaction with an environment, it is currently the most popular benchmark to evaluate different RL fine-tuning algorithms.
@@ 7.1 Task Descriptions
Mathematical reasoning. We evaluate on mathematical problem-solving using the MATH dataset  (hendrycksmath2021), which consists of different competition math problems of varying level of difficulty. A score of $1$ is achieved if the agent solves the problem and outputs an answer that is correct and properly formatted. We evaluate on a subset of $500$ problems from the test dataset of the highest difficulty level, which we call MATH500-Hard. The remaining $12,000$ problems are used as the training set for RL fine-tuning.
Dialogue game. We use the popular game of 20 Questions (20Q) as a representative strategic dialogue task, where the LLM agent acts as the guesser to uncover the hidden word by an oracle. 20Q was chosen because it was non-adversarial (so we can evaluate against a fixed LLM as the oracle), and requires the LLM agent to generate a cohesive sequence of actions over multiple steps. Though many implementations exist  (srivastava2023beyond; abdulhai2023lmrl), we follow the one by 20Q where the set of hidden words can be any in a set of $1,823$ objects from the THINGS dataset (hebart2019things). A reward of $1$ is achieved if the guesser correctly identifies the hidden object within $20$ turns, or questions, where correctness if determined by using the oracle LLM as a judge. We use GPT4.1 (openai2024gpt4technicalreport) as the oracle. We construct a training set of $1,000$ objects and a test set of $500$ different objects through random sampling.
Customer service. We consider $\tau$-bench as a representative LLM agent task that requires a mixture of dialogue and tool-use to solve (yao2024taubenchbenchmarktoolagentuserinteraction). The LLM agent must act as a customer service representative in various scenarios such as modifying items in an user’s order, and follow a rigid set of policy guidelines. At every step, the LLM agent can either communicate with the user, or make an API call that interacts with a backend database. At the end, the agent receives a score of $1$ if the database entries match ground-truth values, and the agent did not violate any policy guidelines via their actions. Users are simulated using a GPT4.1 (openai2024gpt4technicalreport) model prompted with both an initial request (such as modifying or cancelling an order) as well an identity that can be verified using the database. There are two categories of scenarios: (1) in retail, the LLM agent must modify pending orders of items, return or exchange delivered orders, or update user information, and (2) in airline, the LLM agent must book, modify, or cancel flight reservations. To test generalization, we compile a training dataset of $2,500$ user scenarios in the retail category, and evaluate on a test set of $500$ different retail scenarios, as well as $500$ airline scenarios. Note that none of the methods are trained on any airline scenarios.
@@ 7.2 Results
Prompting GPT4.1 ReAct (yao2022react) $\mathbf{95.1}$ $30.2$ $0.44$ $0.32$
Fine-tuning Qwen2.5-7B-Instruct RFT $52.5$ $12.6$ $0.21$ $0.13$
PPO  (schulman2017proximal) $52.3$ $17.2$ $0.28$ $0.14$
GRPO  (shao2024deepseekmathpushinglimitsmathematical) $49.8$ $18.4$ $0.24$ $0.11$
SAC   (ablation) $48.2$ $9.8$ $0.18$ $0.11$
NLRL  (feng2025naturallanguagereinforcementlearning) $62.4$ $25.8$ $0.25$ $0.16$
NLAC (ours) $60.2$ $26.0$ $0.42$ $0.22$
Fine-tuning QwQ-32B RFT $72.5$ $22.0$ $0.35$ $0.29$
PPO  (schulman2017proximal) $71.4$ $24.0$ $0.47$ $0.41$
GRPO  (shao2024deepseekmathpushinglimitsmathematical) $70.8$ $25.6$ $0.48$ $0.39$
SAC   (ablation) $64.7$ $13.2$ $0.31$ $0.21$
NLRL  (feng2025naturallanguagereinforcementlearning) $73.5$ $30.8$ $0.44$ $0.31$
NLAC (ours) $72.7$ $\mathbf{32.1}$ $\mathbf{0.59}$ $\mathbf{0.45}$
@@
We compare NLAC with $k=1$ and $m=1$ against both prompting and fine-tuning baselines. We found those settings of hyperparameters was sufficient to achieve good performance, though more stochastic environments may warrant higher $k$. For baselines that involve fine-tuning, we consider two LLMs: Qwen2.5-7B-Instruct (qwen2.5), and QwQ-32B (qwq32b), which is also trained on reasoning traces. We choose these two LLMs to measure the effect of increasing size and pre-training on reasoning traces on the performance of the RL methods. Our baselines can be categorized into the following (training details can be found in Appendix B.4):
Prompting. We perform ReAct prompting  (yao2022react) of a state-of-the-art frontier model GPT4.1 (openai2024gpt4technicalreport). Because such models do not expose weights for RL fine-tuning, we rely on the zero-shot capabilities of the LLM without any additional training on the tasks.
Rejection fine-tuning. We perform rejection fine-tuning (RFT) where at every iteration, the base LLM policy collects a set of on-policy rollouts. We append only the successful rollouts to a buffer, then train the LLM using SFT on samples from the buffer.
RL fine-tuning. The most standard way to perform RL fine-tuning is to train the LLM to optimize score using a policy gradient algorithm on on-policy rollouts. We consider both PPO  (schulman2017proximal) and GRPO  (shao2024deepseekmathpushinglimitsmathematical) as the algorithm, the difference being that PPO additionally learns a token-level value function on Monte-Carlo rollouts as a baseline to stabilize reward, whereas GRPO computes the average reward across $8$ different rollouts. We found that increasing the number of rollouts for GRPO only harmed performance.
Ablations. We consider an ablation of our approach that is soft actor-critic (SAC) training. Instead of performing policy iteration in language space, SAC simply learns scalar values via a token-level Q-function using traditional Bellman backups, and then performs policy extraction by fitting the policy to the maximum-entropy policy using the learned Q-function  (haarnoja2018soft). Finally, we compare against NLRL (feng2025naturallanguagereinforcementlearning); during training, instead of using the language Bellman backup to train a successor function and refinement to train the policy, NLRL aggregates environment transitions in-context with predicted future value, then enumerates multiple actions and their values, respectively. To keep NLRL tractable, as the state and action space are prohibitively large, we limit to $8$ transitions and actions.
@@
The results of our evaluation are presented in Table 1. We see that for both LLM models, NLAC outperforms other fine-tuning approaches, and even prompting of a significantly stronger LLM, particularly on long-horizon tasks. The only task where NLAC matches other methods is mathematical reasoning, which is a single-step task, in which NLAC reduces to performing self-refinement using a generative reward model (madaan2023selfrefineiterativerefinementselffeedback); this is because in single-step tasks, our natural language critic is only trained to predict reward. Meanwhile, on tasks requiring multi-step interaction, which our method is tailored for, NLAC greatly outperforms all baselines, achieving a $30\%$ improvement in 20Q and $\tau$-retail over standard RL fine-tuning. Furthermore, as shown in Figure 4, NLAC also requires fewer gradient steps to achieve maximum performance, illustrating the effectiveness of leveraging text-based reasoning. Surprisingly, ablation SAC performed worst; this can be attributed to the fact that token-level Q-values are difficult to estimate precisely, so directly fitting the policy to these Q-values can hurt performance. NLRL performed comparably to NLAC on math and 20Q, though it is important to note that NLRL effectively sees $8\times$ more samples during training. However, NLRL performed noticeably worse on $\tau$-bench. Qualitatively, we found that values learned by NLRL were almost always positive, which made them not helpful for policy improvement; this could be due to a combination of poor modeling of environment dynamics due to limited in-context samples, and the implicit bias of instruction fine-tuned models. Overall, this serves as evidence that NLAC is more suitable to complex tasks.
@@
We demonstrate qualitative examples of how NLAC improves the base LLM policy in Figure 2 and Figure 3 for 20Q and $\tau$-bench, respectively. In 20Q, base LLM agents often resort to linear search over some specific characteristic of an object, when it is likely more optimal to further explore over other discriminators. In the example, the critique by our natural language critic explicitly corrects this linear strategy when it occurs. Meanwhile, in $\tau$-bench, one of the most common failure modes is partial resolution of complex requests, especially when the agent must also follow complicated dynamics and rules. In the example, the agent is told that the user wants to make “a couple of exchanges," but according to policy guidelines, modifications to the database can only be done via one tool-call per rollout. Therefore, the agent should not make a tool-call to exchange the first item, but instead collect all items to be exchanged into a single call in the future. This kind of error would be difficult to correct with just a scalar reward as feedback. However, the critique by our natural language critic identifies exactly which policy guideline would be violated, allowing for the LLM agent to easily understand and correct the error.

### v2

@@ 4.2 Policy Improvement
Thus far, we showed how to train the natural language critic to evaluate a fixed policy $\pi$. We now show how an improved policy can be learned using textual critiques $Q_{L}^{\pi}(s,a)$ obtained by a critic using Equation 4.1. Naturally, such policy is a greedy policy where $a\sim\pi(\cdot|s)$ satisfies $a=\arg\max_{a^{\prime}}Q^{\pi}_{L}(s,a^{\prime})$. Note that we assume:
@@
For any policy $\pi$, the set $\{Q_{L}^{\pi}(s,a^{\prime})\}_{a^{\prime}\in\mathcal{A}}$ computed using Equation 4.1 for any state $s$ forms a totally-ordered set with binary relation $\geq$.
@@
However, computing the greedy policy is intractable for LLM agent tasks, where the action spaces $\mathcal{A}\subseteq\mathcal{V}^{*}$ are combinatorial in the token vocabulary, making it impossible to enumerate all possible actions to find the optimal one. While prior works have proposed sampling a subset of actions and reweighting (Li et al., 2024), we find empirically that for tractable sample sizes, this approach does not sufficiently explore the space of possible actions.
@@
Finally, we propose a policy improvement objective from $\pi$ to $\pi^{\prime}$ that projects the refinement policy back to the base policy, similar to the policy updates in SAC (Haarnoja et al., 2018). However, rather than parameterizing a target policy using the learned values, which requires enumeration over actions and is intractable in our setting, we let the target policy be the refinement policy:
$\pi^{\prime}=$ (5)
$\arg\max_{\pi}\;\mathbb{E}_{s_{t}\sim\mathcal{D}}\Bigl[D_{f}\bigl(\pi(\cdot|s_{t})\|\pi^{r}(\cdot|\ldots,Q^{\pi}_{L}(s_{t},a_{t}^{m}))\bigr)\Bigr]\,.$
@@
4.3 Theoretical Results
Due to space, we provide a theoretical analysis of NLAC in Appendix A. At a high level, our analysis connects the representations used by our learned critic to successor features, which allows us to recover standard scalar value-learning guarantees even though the critic outputs text.
Our main results state that, under mild realizability assumptions on the reward and on how the language Bellman backup composes representations, we show that the learned language critic can be used to recover the true Q-function:
@@ 4.3 Theoretical Results
Consider policy evaluation via Equation 4.1 and let $Q_{L}^{\pi}$ be the natural language critic at convergence. For any state $s$ and action $a$, there exists monotonic mapping $g$ such that $Q^{\pi}(s,a)=g(Q_{L}^{\pi}(s,a))$, where $Q^{\pi}$ denotes the true scalar Q-function.
@@ 5.1 Training Components
Policy. Many prior works have parameterized policies as LLMs. One of the greatest advantages of doing so is the ability to leverage the strong reasoning capabilities of LLMs from chain-of-thought prompting (Wei et al., 2023; Yao et al., 2022). By choosing a proper prompt $p_{\text{react}}$, an LLM policy can be instructed to describe their underlying thoughts for choosing a particular action in addition to generating the action itself $a_{t}\sim\mathcal{M}_{\theta}(p_{\text{react}}(s_{t}))$.
Language successor model. LLMs have demonstrated efficacy at predicting realistic future rollouts in a variety of environments (Lin et al., 2024). These futures are generated by simply processing the state-action in a prediction prompt $p_{\text{pred}}$ that also instructs the LLM to summarize rollouts into concise textual descriptions, then sampling from the LLM output $M_{\theta}(\cdot\mid s_{t},a_{t})=\mathcal{M}_{\theta}(p_{\text{pred}}(s_{t},a_{t}))$.
@@
$\mathcal{B}_{L}M_{\theta}(\cdot\mid s_{t},a_{t})=\mathcal{M}_{\theta}\bigl(p_{\text{tpred}}(r_{t},s_{t+1},d_{t+1})\bigr)\,,\ \text{where}$
$\quad d_{t+1}\sim\mathcal{M}_{\theta}\bigl(p_{\text{pred}}(s_{t+1})\bigr)\,.$
@@
$Q_{L}^{\theta}(s_{t},a_{t})\sim\mathcal{M}_{\theta}(p_{\text{eval}}(d_{t}^{(1)},\ldots,d_{t}^{(k)}))\,.$
Refinement policy. Finally, the refinement policy $\pi^{r}$ can also be obtained by an LLM instructed to refine its latest action given an evaluation similar to prior self-refinement approaches (Madaan et al., 2023). The refined action is obtained via prompt $p_{\text{refine}}$ as
$a_{t}^{r}\sim\mathcal{M}_{\theta}(p_{\text{refine}}(s_{t},a_{t}^{1},\ldots,Q^{\theta}_{L}(s_{t},a_{t}^{m})))\,.$
@@ 5.2 Training Algorithm
$\mathcal{L}_{1}(s_{t},a_{t},s_{t+1})=D_{\mathsf{KL}}\left(\mathcal{B}_{L}M_{\bar{\theta}}(\cdot|s_{t},a_{t})\|M_{\theta}(\cdot|s_{t},a_{t})\right)\,$,
where $\bar{\theta}$ are reference parameters that are an exponentially moving average of the trained parameters, in order to prevent generative collapse  (Shumailov et al., 2024). We choose the reverse direction of KL-divergence to capture the full diversity over possible futures. Then, for policy improvement, we train the policy on the log-likelihood loss:
$\mathcal{L}_{2}(s_{t},k)=-\log\pi_{\theta}(a_{t}^{r}|s_{t})\,,\ \text{where}$
$\quad a_{t}\!\sim\!\pi_{\theta}(\cdot\mid s_{t})\,,\ a_{t}^{r}\!\sim\!\pi^{r}_{\theta}(\cdot|\ldots,Q_{L}^{\theta}(s_{t},a_{t}))\,.$
This objective can be interpreted as distillation, but using generations by the refinement policy as the teacher policy. Note that the loss depends on $k$ via $Q_{L}^{\theta}(s_{t},a_{t})$ given by Equation 4.1. Note that by default, our refinement policy relies on the base reasoning capabilities of the pretrained LLM. In Appendix B.5, we show results when the refinement policy is explicitly trained.
@@
We show pseudocode for NLAC in Algorithm 1. Like other methods that utilize pretrained LLMs, our method is susceptible to catastrophic forgetting. We were able to avoid this in our experiments by training in low-data regimes. However, we discuss effective methods for mitigating catastrophic forgetting in Appendix B.6.
@@ 6 Experiments
To demonstrate the effectiveness of NLAC, we evaluate our method on a variety of LLM agent tasks: mathematical reasoning (Hendrycks et al., 2021), strategic dialogue  (Pu, 2023), and customer service using mixed dialogue and tool-use (Yao et al., 2024). Though mathematical reasoning does not involve interaction with an environment, it is currently the most popular benchmark to evaluate different RL fine-tuning algorithms.
@@ 6.1 Task Descriptions
Mathematical reasoning. We evaluate on mathematical problem-solving using the MATH dataset  (Hendrycks et al., 2021), which consists of different competition math problems of varying level of difficulty. A score of $1$ is achieved if the agent solves the problem and outputs an answer that is correct and properly formatted. We evaluate on a subset of $500$ problems from the test dataset of the highest difficulty level, which we call MATH500-Hard. The remaining $12,000$ problems are used as the training set for RL fine-tuning.
Dialogue game. We use the popular game of 20 Questions (20Q) as a representative strategic dialogue task, where the LLM agent acts as the guesser to uncover the hidden word by an oracle. 20Q was chosen because it was non-adversarial (so we can evaluate against a fixed LLM as the oracle), and requires the LLM agent to generate a cohesive sequence of actions over multiple steps. Though many implementations exist  (Srivastava et al., 2023; Abdulhai et al., 2023), we follow the one by Pu (2023) where the set of hidden words can be any in a set of $1,823$ objects from the THINGS dataset (Hebart et al., 2019). A reward of $1$ is achieved if the guesser correctly identifies the hidden object within $20$ turns, or questions, where correctness if determined by using the oracle LLM as a judge. We use GPT5 (Singh et al., 2025) as the oracle. We construct a training set of $1,000$ objects and a test set of $500$ different objects through random sampling.
Customer service. We consider $\tau$-bench as a representative LLM agent task that requires a mixture of dialogue and tool-use to solve (Yao et al., 2024). The LLM agent must act as a customer service representative in various scenarios such as modifying items in an user’s order, and follow a rigid set of policy guidelines. At every step, the LLM agent can either communicate with the user, or make an API call that interacts with a backend database. At the end, the agent receives a score of $1$ if the database entries match ground-truth values, and the agent did not violate any policy guidelines via their actions. Users are simulated using a GPT5 (Singh et al., 2025) model prompted with both an initial request (such as modifying or cancelling an order) as well an identity that can be verified using the database. There are two categories of scenarios: (1) in retail, the LLM agent must modify pending orders of items, return or exchange delivered orders, or update user information, and (2) in airline, the LLM agent must book, modify, or cancel flight reservations. To test generalization, we compile a training dataset of $2,500$ user scenarios in the retail category, and evaluate on a test set of $500$ different retail scenarios, as well as $500$ airline scenarios. Note that none of the methods are trained on any airline scenarios.
@@ 6.2 Results
Prompting GPT5 ReAct (Yao et al., 2022) $80.8$ $30.2$ $0.44$ $0.32$
Fine-tuning Qwen2.5-7B-Instruct RFT $49.8$ $12.6$ $0.21$ $0.13$
PPO  (Schulman et al., 2017) $52.3$ $17.2$ $0.31$ $0.14$
GRPO  (Shao et al., 2024) $52.5$ $18.4$ $0.32$ $0.11$
NLRL  (Feng et al., 2025)) $53.4$ $25.8$ $0.25$ $0.16$
Self-Distillation (ablation) $\mathbf{57.1}$ $12.4$ $0.24$ $0.14$
NLQL (ablation) $56.4$ $24.2$ $0.29$ $0.19$
NLAC (ours) $56.2$ $\mathbf{26.0}$ $\mathbf{0.42}$ $\mathbf{0.22}$
Fine-tuning QwQ-32B RFT $69.4$ $22.0$ $0.35$ $0.29$
PPO  (Schulman et al., 2017) $\mathbf{73.4}$ $24.0$ $0.52$ $0.41$
GRPO  (Shao et al., 2024) $71.8$ $25.6$ $0.49$ $0.39$
NLRL  (Feng et al., 2025)) $71.5$ $31.8$ $0.44$ $0.31$
Self-Distillation (ablation) $70.1$ $26.8$ $0.35$ $0.29$
NLQL (ablation) $71.2$ $31.4$ $0.51$ $0.36$
NLAC (ours) $72.7$ $\mathbf{32.1}$ $\mathbf{0.56}$ $\mathbf{0.43}$
@@
We compare NLAC with $k=1$ and $m=1$ against both prompting and fine-tuning baselines. We found those settings of hyperparameters was sufficient to achieve good performance, though more stochastic environments may warrant higher $k$. For baselines that involve fine-tuning, we consider two LLMs: Qwen2.5-7B-Instruct (Yang et al., 2024), and QwQ-32B (Team, 2025), which is also trained on reasoning traces. We choose these two LLMs to measure the effect of increasing size and pre-training on reasoning traces on the performance of the RL methods. Our baselines can be categorized into the following (training details can be found in Appendix B.4):
Prompting. We perform ReAct prompting  (Yao et al., 2022) of a state-of-the-art frontier model GPT5 (OpenAI et al., 2024). Because such models do not expose weights for RL fine-tuning, we rely on the zero-shot capabilities of the LLM without any additional training on the tasks.
Rejection fine-tuning. We perform rejection fine-tuning (RFT) where at every iteration, the base LLM policy collects a set of on-policy rollouts. Then, we train the LLM using SFT only on the successful rollouts.
RL fine-tuning. The most standard way to perform RL fine-tuning is to train the LLM to optimize score using a policy gradient algorithm on on-policy rollouts. We consider both PPO  (Schulman et al., 2017) and GRPO  (Shao et al., 2024) as the algorithm, the difference being that PPO additionally learns a token-level value function on Monte-Carlo rollouts as a baseline to stabilize reward, whereas GRPO computes the average reward across $8$ different rollouts. We found that increasing the number of rollouts harmed performance. Finally, we compare against NLRL (Feng et al., 2025). Instead of training a successor function, NLRL aggregates on-policy rollouts in-context; and intead of distilling from a refinement policy, NLRL enumerates multiple actions and distills the action with highest language value. To keep NLRL tractable, as the rollout and action space are prohibitively large to fit in-context, we condition the critic and policy on $4$ rollouts and actions in-context.
Ablations. The main algorithmic contributions of NLAC involve learning a language successor function to tractably extract values in language space, then distilling from a refinement policy that conditions on the learned values. We consider two separate ablations that separately remove each contribution. First, we compare against Self-Distillation, where instead of learning and conditioning on language values, our refinement policy conditions on on-policy rollouts, similar to Shinn et al. (2023); again, for tractability, we fit $4$ such rollouts in-context. Finally, we compare against NLQL (Q-learning instead of actor-critic), where instead of distilling from a refinement policy, we sample $4$ actions and distil the one that achieves the highest value (by sentiment). NLQL can also be considered an instance of NLRL due also employing value learning in language space.
@@
The results of our evaluation are presented in Table 1. We see that for both LLM models, NLAC outperforms other fine-tuning approaches, and even prompting of a significantly stronger LLM, particularly on long-horizon tasks. The only task where NLAC matches other methods is mathematical reasoning, which is a single-step task, in which NLAC reduces to performing self-distillation using a generative reward model (Madaan et al., 2023); this is because in single-step tasks, our natural language critic is only trained to predict reward. Meanwhile, on tasks requiring multi-step interaction, which our method is tailored for, NLAC greatly outperforms all baselines, achieving a $30\%$ improvement in 20Q and $\tau$-retail over standard RL fine-tuning. Furthermore, as shown in Figure 4, NLAC also requires fewer gradient steps to achieve maximum performance, illustrating the effectiveness of leveraging text-based reasoning. NLRL performed comparably to NLAC on math and 20Q, though it is important to note that NLRL effectively sees $4\times$ more samples during training. However, NLRL performed noticeably worse on $\tau$-bench. Qualitatively, we found that values learned by NLRL were almost always positive, which made them not helpful for policy improvement; this could be due to a combination of poor modeling of environment dynamics due to limited in-context samples, and the implicit sycophancy of instruction fine-tuned models. Finally, from our ablation analysis, we found that both novel components were important. Self-Distillation proved ineffective on tasks with long, complex rollouts, showing the importance of learning compact summarizations of future. NLQL performed better, but we found that sampling multiple actions explored the action space much more poorly than a refinement policy that uses critic output. Overall, this serves as evidence that NLAC is more suitable to complex tasks.
@@
In Figure 2, base LLMs agents in 20Q often resort to linear search over some specific characteristic of an object, when it is likely more optimal to further explore over other discriminators. In the example, the critique by our natural language critic explicitly corrects this linear strategy when it occurs. Meanwhile, as shown in Figure 3, one of the most common failure modes in $\tau$-bench is the partial resolution of complex requests, especially when the agent must also follow complicated dynamics and rules. In the example, the agent is told that the user wants to make “a couple of exchanges," but according to policy guidelines, modifications to the database can only be done via one tool-call per rollout. Therefore, the agent should not make a tool-call to exchange the first item, but instead collect all items to be exchanged into a single call in the future. This kind of error would be difficult to correct with just a scalar reward as feedback. However, the critique by our natural language critic identifies exactly which policy guideline would be violated, allowing for the LLM agent to easily understand and correct the error.